A 20 meter wooden pole is cut into two parts that have the ratio 3 to 2. How much longer is the larger part than the smaller?
Answer by anshu_a(16)

lenth of the pole = 20m
ratio of parts = 3:2
large part = 20*(3/5) = 12
small part = 20-12 = 8
lenth of large part more than smaller part = 12-8 = 4m